MATEMATICA Anno Accademico 2005-2006

Scienze Geologiche M.Novaga, V.M. Tortorelli

————————————————————————————————————————–

I prova finale, 2 febbraio 2006

RISULTATI

————————————————————————————————————————–

PIERETTI Carlo  sufficiente